OpenStudy (anonymous):

Please Explain How To Do This Step By Step Im So Stuck. Thank You 6x-4y=54

OpenStudy (anonymous):

Are you solving for x or y?

OpenStudy (anonymous):

Im not sure it dosent say in my book just solve and show the steps thats all

OpenStudy (anonymous):

Okay well then lets just do both. You sure this is all the information? There isn't a second equation?

OpenStudy (anonymous):

lol Yes there is right Under it Hold on one Sec

OpenStudy (anonymous):

-9x+2y=-69

OpenStudy (anonymous):

Okay so we have: 6x-4y=54 -9x +2y = -69

OpenStudy (anonymous):

Yes Thats It :)

OpenStudy (nincompoop):